OpenStudy (anonymous):

construct a box plot for these data and identify any outliers : 25 22 26 23 27 26 28 18 25 24 12 . this is statistics course if you know can you show me thx

OpenStudy (anonymous):

put the numbers in numerical order... 12, 18, 22, 23, 24, 25, 25, 26, 26, 27, 28 the median (middle number) also known as Q2 = 25 Q1 is the middle number of all the numbers before the median : 12, 18, 22, 23, 24 , so Q1 = 22 Q3 is the middle number of all the numbers after the median : 25, 26, 26, 27, 28, so Q3 = 26 Lower extreme (lowest number) is 12 Upper extreme (highest number) is 28 Interquartile range is the length of the box. An outlier is any value that lies more then 1 1/2 times the length of the box from either end of the box. If the data point is below Q1 - 1.5 x IQR or above Q3 + 1.5 x IQR. The IQR is Q3 - Q1, which is 26 - 22 = 4 Outliers would be : Q1 - 1.5(4) Q3 + 1.5(4) 22 - 1.5(4) 26 + 1.5(4) 22 - 6 = 16 26 + 6 = 32 so any number below 16 or above 32 are your outliers. Therefore, your outiers are 12.

OpenStudy (anonymous):

prefect thank so much. my teacher show more complicate

OpenStudy (anonymous):

you already take statics or you take it now? do you often get online? i really need have with statistics.

OpenStudy (anonymous):

 I can't draw very well...lol..